OpenStudy (anonymous):

2. Evaluate 5m – 2n for m = –3 and n = –2. (1 point) –19 –11 19 11

OpenStudy (anonymous):

5*-3=-15 2*-2=-4 -15--4=-11

OpenStudy (anonymous):

3. Simplify (4x + 1) + (–3x + 9) (1 point) x + 10 7x – 8 7x + 10 x – 8

OpenStudy (anonymous):

That one is pretty straight forward as well. + before a ( ) mean that nothing withing ( ) changes, so it's 4x + 1 -3x +9 = answer

OpenStudy (anonymous):

so it would be x + 10 ?

OpenStudy (anonymous):

yes